# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ 2.1 High Dimensional Predictive Regression
Suppose that a time series of outcome $y_{t}$ is generated by the following linear predictive regression: [^1] For simplicity of exposition, an intercept in (2.1) is omitted, without loss of generality, as explained by Mei and Shi, (2024) that in LASSO the intercept can be handled by the well-known Frisch-Waugh-Lovell theorem. In practical implementation — throughout all simulations and empirical exercises in this paper — we keep an unpenalized intercept in the model.
@@
where the error term $u_{t}$ is a stationary and martingale difference sequence (m.d.s.) with mean zero and conditional variance $\sigma_{u}^{2}$. We consider two types of regressors with different stochastic properties: $X_{t}=(x_{1,t},\cdots,x_{p_{x},t})^{\top}$ is a $p_{x}\times 1$ vector of unit roots satisfying $X_{t}=X_{t-1}+e_{t}$ with $e_{t}=(e_{1,t},\cdots,e_{p_{x},t})^{\top}$ being a $p$-dimensional vector of stationary time series, and $Z_{t}=(z_{1,t},\cdots,z_{p_{z},t})^{\top}$ is a $p_{z}\times 1$ vector of stationary regressors. The two types of regressors are combined into a long vector $W_{t}=(X_{t}^{\top},Z_{t}^{\top})^{\top}=(w_{1,t},\cdots,w_{p,t})^{\top}$ of $p$ ($=p_{x}+p_{z}$) elements, and the associated coefficients are placed into $\theta^{*}=(\beta^{*\top},\gamma^{*\top})^{\top}\in\mathbb{R}^{p}$. For simplicity, let the initial value $\|W_{t=0}\|_{\infty}=O_{p}(1)$. Define the sparsity index $s=\sum_{j=1}^{p}\boldsymbol{1}\left\{\theta_{j}^{*}\neq 0\right\}$ as the number of nonzero components in the vector of coefficients $\theta^{*}$.
As in the default R program option glmnet::glmnet(x,y), it is a common practice in LASSO to scale-standardize each regressor $w_{j,t}$ by its sample standard deviation (s.d.)  $\widehat{\sigma}_{j}=\sqrt{\frac{1}{n}\sum_{t=1}^{n}(w_{j,t-1}-\bar{w}_{j})^{2}},$ where $\bar{w}_{j}=n^{-1}\sum_{t=1}^{n}w_{j,t-1}$ is the sample mean. The diagonal matrix $D={\rm diag}\left(\widehat{\sigma}_{1},\widehat{\sigma}_{2},\cdots,\widehat{\sigma}_{p}\right)$ stores the sample standard deviations. The standardized LASSO (Slasso) estimator is
@@
Compared to the plain LASSO estimator that replaces $D$ by the identity matrix, the Slasso estimator is scale-invariant: if the regressor $w_{j,t-1}$ is multiplied by a nonzero constant $m$, then the $j$-th coefficient estimator changes proportionally into $\widehat{\theta}_{j}^{{\rm S}}/m.$ The standardization renders the magnitudes of unit roots into the same order as those of stationary regressors, so that the same LASSO tuning parameter $\lambda$ in (2.2) is valid for both stationary and persistent regressors.
@@ 2.2 Two Types of Biases
When data are i.i.d. or stationary, LASSO is subject to shrinkage bias as well as nonstandard asymptotic distribution which cannot be used for standard inference (Fu and Knight,, 2000). This motivates Zhang and Zhang, (2014) to bring forth the desparsified LASSO to correct the bias and recover asymptotic normality. With high dimensional unit roots, $\widehat{\theta}^{{\rm S}}$ is subject to the shrinkage bias, while the highly persistent unit root regressors incur the Stambaugh bias, further distorting the standard t-statistics inference. In this section, we examine both the shrinkage bias and the Stambaugh bias, and propose XDlasso for correcting both biases.
We are interested in inference for a null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$ for some $j\in[p]$, a prevalent practice in empirical studies. [^2] A joint inference for the fixed-dimensional subvector $\theta_{\mathcal{J}}^{*}=(\theta_{j}^{*})_{j\in\mathcal{J}}$ is feasible for a subset $\mathcal{J}\subset[p]$. However, a subvector will incur additional complexity when persistent and stationary time series are mixed in the high dimensional regressors. For simplicity of exposition, we focus on inference for a scalar coefficient. In a low dimensional linear regression where $p$ is fixed, the Frisch-Waugh-Lovell theorem yields the following formulation of the ordinary least squares (OLS) estimator:
@@
where $w_{j}^{\perp}=(w_{j,0}^{\perp},\dots,w_{j,n-1}^{\perp})^{\top}$ is the OLS residual from regressing $w_{j,t}$ on all other regressors $W_{-j,t}=(w_{k,t})_{k\neq j}$. OLS induces a large variance as $p$ gets large, and becomes infeasible when $p>n$. Now, consider replacing the OLS residual $w_{j}^{\perp}$ by a generic score vector $r_{j}=(r_{j,0},\dots,r_{j,n-1})^{\top}$ to construct an estimator of $\theta_{j}^{*}$ that is linear in $y_{t}$ in the form of $\widehat{\theta}_{j}^{({\rm lin})}=\frac{\sum_{t=1}^{n}r_{j,t-1}y_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}.$ Since $y_{t}=w_{j,t-1}\theta_{j}^{*}+W_{-j,t-1}^{\top}\theta_{-j}^{*}+u_{t}$ where $\theta_{-j}^{*}=(\theta_{k}^{*})_{k\neq j}$ is the vector of coefficients excluding the $j$-th entry, the generic estimator $\widehat{\theta}_{j}^{({\rm lin})}$ can be decomposed into
$\widehat{\theta}_{j}^{({\rm lin})}=\theta_{j}^{*}+\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}\theta_{-j}^{*}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}.=:\theta_{j}^{*}+{\rm N}_{j}+{\rm B}_{j},$
@@
For OLS, the score vector $r_{j}=w_{j}^{\perp}$ is orthogonal to the column space of $W_{-j,\cdot}:=(W_{-j,0},\dots,W_{-j,n-1})^{\top}$, under which ${\rm B}_{j}=0$ and no bias is present. The bias term ${\rm B}_{j}$ pops up whenever $r_{j}$ is not orthogonal to $W_{-j,\cdot}$, which happens if we add a penalty to the OLS objective function. With LASSO penalty we call ${\rm B}_{j}$ shrinkage bias.
Following Zhang and Zhang, (2014), we replace the unknown parameter $\theta_{-j}^{*}$ by the feasible workhorse estimator $\widehat{\theta}_{-j}^{{\rm S}}=(\widehat{\theta}_{k}^{{\rm S}})_{k\neq j}$:
@@
where $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ is the Slasso residual. Though LASSO may shrink $\widehat{\theta}_{j}^{{\rm S}}$ all the way to exactly zero, the second term in (2.4) is continuously distributed and therefore $\widehat{\theta}_{j}$ is a desparsified version of $\widehat{\theta}_{j}^{{\rm S}}$. Elementary calculation yields
$\widehat{\theta}_{j}-\theta_{j}^{*}=\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j}^{{\rm S}})}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}=:{\rm N}_{j}-(\widehat{{\rm B}}_{j}-{\rm B}_{j}),$ (2.5)
@@
Let $\omega_{j}$ denote the standard deviation of ${\rm N}_{j}$. To secure asymptotic normality for ${\rm N}_{j}$, we need an appropriate score vector $r_{j}$ such that as $n\to\infty$:
@@
The result (R1) is well established by Zhang and Zhang, (2014) under i.i.d. data when the score $r_{j,t}$ is taken as the LASSO residual of regressing $w_{j,t}$ on all other regressors. However, when $w_{j,t}$ is a unit root, this practice leads to a highly persistent $r_{j,t}$ and ruins the asymptotic normality of ${\rm N}_{j}$. The issue that highly persistent regressors produce an asymptotically non-normally distributed OLS estimator that is skewed away from zero is called the Stambaugh bias (Stambaugh,, 1999). To safeguard (R1), we must seek a score that is less persistent than a unit root.
@@
${\rm\widehat{B}}_{j}-{\rm B}_{j}$ $=-g_{j}^{\top}(\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}),\ \text{where}\ g_{j}=\dfrac{\sum_{t=1}^{n}W_{-j,t-1}r_{j,t-1}}{\sum_{t=1}^{n}w_{j,t-1}r_{j,t-1}}\cdot$ (2.6)
Note that $|{\rm\widehat{B}}_{j}-{\rm B}_{j}|\leq\|g_{j}\|_{\infty}\cdot\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}.$ Slasso’s $L_{1}$ estimation error $\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}$ is invariant with the choice of score vector $r_{j,t}$, and diminishes if $\widehat{\theta}^{{\rm S}}$ is consistent. Thus, it suffices to control the order of $\|g_{j}\|_{\infty}$ to achieve (R2). The explicit expression of $g_{j}$ in (2.6) suggests that a weak correlation in sup-norm between $W_{-j,t}$ and the score $r_{j,t}$ (relative to the correlation between $w_{j,t}$ and $r_{j,t}$) helps.
With these routes in mind, we devise XDlasso in the following section. It proceeds with two key steps: (i) conducting an IVX transformation to get a new variable less persistent than the unit root regressors to eliminate the Stambaugh bias; (ii) running an auxiliary LASSO regression to construct an $r_{j,t}$ to remove the shrinkage bias.
@@ 2.3 IVX-Desparsified LASSO
For inference of $\theta_{j}^{*}$ in low dimensions, Phillips and Magdalinos, (2009)’s IVX method adopts a generated instrument by $w_{j,t}$. When $w_{j,t}$ has a unit root, this self-generated instrument is mildly integrated. The mitigation of persistence will remove the Stambaugh bias when the sample size passes to infinity, and thus recovers asymptotic normality of the IVX estimator. Due to the coexistence of Stambaugh bias and shrinkage bias, the wisdom of IVX cannot be directly transplanted into the high dimensional case. Instead, we integrate the IVX method with the idea of desparsified LASSO after comprehending the mechanism of the biases illustrated in Section 2.2.
@@
where $\rho_{\zeta}\in(0,1)$ is a user-determined tuning parameter. Define the sample s.d. of the instrument as $\widetilde{\varsigma}_{j}=\sqrt{n^{-1}\sum_{t=0}^{n-1}\left(\zeta_{j,t}-\bar{\zeta}_{j}\right)^{2}.}$ We unify the scale for unit roots and stationary regressors by standardizing the instrument with its s.d.
$\widetilde{\zeta}_{j,t}=\zeta_{j,t}/\widetilde{\varsigma}_{j}.$ (2.8)
@@
For low dimensional predictive regressions, the IVX literature has established asymptotic normality of ${\rm N}_{j}$ in (2.5) taking $r_{j,t}=\widetilde{\zeta}_{j,t}$; see Phillips and Magdalinos, (2009). When $w_{j,t}$ is a unit root, the IV $\zeta_{j,t}$ is mildly integrated and less persistent than a unit root. Furthermore, recall that the vector of regressors $W_{-j,t}$ includes either unit roots or stationary regressors. Due to different degrees of persistence, the mildly integrated IV $\widetilde{\zeta}_{j,t}$ and the regressors $W_{-j,t}$ are asymptotically uncorrelated. Thus, when $w_{j,t}$ has a unit root, we can choose the score vector as $r_{j,t}=\widetilde{\zeta}_{j,t}$ to deliver a small-order $\|g_{j}\|_{\infty}$ where $g_{j}$ is defined in (2.6). If $w_{j,t}$ is stationary, however, this score vector fails. When $n$ is large, the instrumental variable $\zeta_{j,t}$ behaves similarly as the stationary regressor $w_{j,t}$, and thus its correlation to the high dimensional stationary regressors in $W_{-j,t}$ is not negligible.
@@
The above analysis implies that we must decorrelate the score vector with the other regressors $W_{-j,t}$ to reduce the order of $g_{j}$ to control for the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$. This decorrelation will produce a unified testing approach for both unit roots and stationary regressors. To this end, we construct a “residual” score vector $\widehat{r}_{j}=(\widehat{r}_{j,0},\cdots,\widehat{r}_{j,n-1})^{\top}$ by the following auxiliary LASSO regression
$\widehat{r}_{j,t}$ $=\widetilde{\zeta}_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ (2.9)
$\widehat{\varphi}^{(j)}$ $=\arg\min_{\varphi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{\zeta}_{j,t-1}-W_{-j,t-1}^{\top}\varphi)^{2}+\mu\|D_{-j}\varphi\|_{1}$ (2.10)
@@
where $\mu$ is the LASSO tuning parameter, and $D_{-j}={\rm diag}(\{\widehat{\sigma}_{k}\}_{k\neq j})$. In low dimensional multivariate predictive regressions, IVX transforms each regressor into a less persistent instrumental variable parallel to (2.7), and constructs a two-stage least squares estimator using all the self-generated instrumental variables. In contrast, we only transform the variable of interest $w_{j,t}$ and estimate one auxiliary regression (2.10).
@@
The score vector $\widehat{r}_{j}$ by (2.9) accommodates stationary and unit root regressors. When $w_{j,t}$ is stationary, the magnitude of the instrument $\widetilde{\zeta}{}_{j,t}$ shares the same order as that of $w_{j,t}$. Thus, the score $\widehat{r}_{j}$ is asymptotically equivalent to the residual of the LASSO regression of $w_{j,t}$ on $W_{-j,t}$. The latter is exactly the score in Zhang and Zhang, (2014) for cross-sectional data, which is also used in Adamek et al., (2023) for stationary time series. When $w_{j,t}$ has a unit root, the instrument $\zeta_{j,t}$ is mildly integrated and has a different degree of persistence from $W_{-j,t}$. Therefore, $\zeta_{j,t}$ is asymptotically uncorrelated to each regressor in $W_{-j,t}$. For notational conciseness, define $\widetilde{W}_{-j,t}=(D_{-j})^{-1}W_{-j,t}$ and $\widetilde{\varphi}^{(j)}=D_{-j}\widehat{\varphi}^{(j)}$. The analysis above suggests $\widetilde{\varphi}^{(j)}\approx 0$ so that $\widehat{r}_{j,t}=\widetilde{\zeta}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\varphi}^{(j)}\approx\widetilde{\zeta}_{j,t}$. Recall from the discussion below (2.8) that the standardized instrument $\widetilde{\zeta}_{j,t}$ is a valid score process for a unit root regressor, and thus the score $\widehat{r}_{j,t}$ can also remove the biases asymptotically. As a result, the residual of the auxiliary LASSO regression provides a unified construction of the score for both stationary and nonstationary $w_{j,t}$. It allows practitioners to conduct hypothesis testing on the coefficient in high dimensional predictive regression regardless of the order of integration of $w_{j,t}$. We maintain an agnostic attitude about the persistence of the regressors; in practical implementation, we need not distinguish these two types.

### v3

@@ 2.1 High Dimensional Predictive Regression
Suppose that a time series of the outcome $y_{t}$ is generated by the following linear predictive regression: [^1] For simplicity of exposition, an intercept in (2.1) is omitted, without loss of generality. As explained by MS24, the intercept in LASSO can be handled by the well-known Frisch-Waugh-Lovell theorem. In practical implementation — throughout all simulations and empirical exercises in this paper — we keep an unpenalized intercept in the model.
@@
where the error term $u_{t}$ is a stationary martingale difference sequence (m.d.s.) with mean zero and conditional variance $\sigma_{u}^{2}$. We consider two types of regressors with different stochastic properties. Firstly, the $p_{x}\times 1$ vector $X_{t}=(x_{1,t},\cdots,x_{p_{x},t})^{\top}$ collects the LURs:
$x_{j,t}=\rho_{j}^{*}x_{j,t-1}+e_{j,t}\text{ for }j=1,2,...,p_{x},$ (2.2)
where $e_{t}=(e_{1,t},\cdots,e_{p_{x},t})^{\top}$ is a $p_{x}$-dimensional vector of stationary time series. The AR(1) coefficient $\rho_{j}^{*}$ in (2.2) is close to 1 when the sample size $n$ is large, specified as
$\rho_{j}^{*}=1+\frac{c_{j}^{*}}{n}\text{ for }j=1,2,...,p_{x},$ (2.3)
where $c_{j}^{*}\in\mathbb{R}$ is allowed to be negative, positive, or zero. Therefore, our framework accommodates nonstationary regressors that are locally integrated ($c_{j}^{*}<0$), unit roots ($c_{j}^{*}=0$), and locally explosive ($c_{j}^{*}>0$). Secondly, stationary regressors are stored in the $p_{z}\times 1$ vector $Z_{t}=(z_{1,t},\cdots,z_{p_{z},t})^{\top}$. The two types of regressors are combined into a long vector $W_{t}=(X_{t}^{\top},Z_{t}^{\top})^{\top}=(w_{1,t},\cdots,w_{p,t})^{\top}$ of $p$ ($p_{x}+p_{z}$) elements, and the associated coefficients are placed into $\theta^{*}=(\beta^{*\top},\gamma^{*\top})^{\top}\in\mathbb{R}^{p}$. Following the literature, we refer to $W_{t}$, which has multiple degrees of persistence, as mixed root regressors. For simplicity, let the initial value $\|W_{t=0}\|_{\infty}=O_{p}(1)$. Define the sparsity index $s=\sum_{j=1}^{p}\boldsymbol{1}\{\theta_{j}^{*}\neq 0\}$ as the number of nonzero components in the coefficient vector $\theta^{*}$.
As in the default R program option glmnet::glmnet(x,y), it is a common practice in LASSO to scale-standardize each regressor $w_{j,t}$ by its sample standard deviation (s.d.) $\widehat{\sigma}_{j}=\sqrt{\frac{1}{n}\sum_{t=1}^{n}(w_{j,t-1}-\bar{w}_{j})^{2}},$ where $\bar{w}_{j}=n^{-1}\sum_{t=1}^{n}w_{j,t-1}$ is the sample mean. Let the diagonal matrix $D={\rm diag}\left(\widehat{\sigma}_{1},\widehat{\sigma}_{2},\cdots,\widehat{\sigma}_{p}\right)$ store the sample standard deviations. The standardized LASSO (Slasso) estimator is
@@
The Slasso estimator is scale-invariant: if the regressor $w_{j,t-1}$ is multiplied by a nonzero constant $m$, then the $j$-th coefficient estimator changes proportionally to $\widehat{\theta}_{j}^{{\rm S}}/m.$ The standardization renders the magnitudes of LURs into the same order as those of stationary regressors, so that the same LASSO tuning parameter $\lambda$ in (2.4) is valid for both stationary and persistent regressors. In contrast, the plain LASSO (Plasso) with the matrix $D$ in (2.4) replaced with the identity matrix is scale-variant. What is worse, equipped with a single tuning parameter $\lambda$, Plasso favors the LURs of a larger order, and shrinks the coefficients of stationary regressors with a smaller order all the way to zero — thus becoming inconsistent.
Remark 1.
We follow the default option of the statistical software to use the s.d. for scaling in the Slasso estimator (2.4). The purpose of scaling is to ensure consistency of Slasso (2.4). The long-run variance, which requires tuning a bandwidth in its estimation, provides no additional benefit. We therefore prefer and stick to the vanilla s.d.
@@ 2.2 Two Types of Biases
When data are i.i.d. or stationary, LASSO is subject to shrinkage bias as well as a nonstandard asymptotic distribution, which cannot be used for standard inference (fu2000asymptotics). This motivates zhang2014confidence to bring forth the desparsified LASSO to correct the bias and recover asymptotic normality. With high dimensional LURs, $\widehat{\theta}^{{\rm S}}$ is subject to not only the shrinkage bias, but also the Stambaugh bias due to persistence, which further distorts the standard t-statistics inference. In this section, we examine both the shrinkage bias and the Stambaugh bias, and propose XDlasso for correcting both biases.
We are interested in inference on a null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$ for a $j\in[p]$, a prevalent practice in empirical studies. In a low dimensional linear regression where $p$ is fixed, the Frisch-Waugh-Lovell theorem yields the following formulation of the ordinary least squares (OLS) estimator:
@@
where $w_{j}^{\perp}=(w_{j,0}^{\perp},\dots,w_{j,n-1}^{\perp})^{\top}$ is the OLS residual from regressing $w_{j,t}$ on all other regressors $W_{-j,t}=(w_{k,t})_{k\neq j}$. OLS induces a large variance as $p$ gets large, and becomes infeasible when $p>n$. Now, consider replacing the OLS residual $w_{j}^{\perp}$ by a generic score vector $r_{j}=(r_{j,0},\dots,r_{j,n-1})^{\top}$ to construct an estimator of $\theta_{j}^{*}$ that is linear in $y_{t}$ in the form of $\widehat{\theta}_{j}^{({\rm lin})}=\frac{\sum_{t=1}^{n}r_{j,t-1}y_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}.$ Since
$y_{t}=w_{j,t-1}\theta_{j}^{*}+W_{-j,t-1}^{\top}\theta_{-j}^{*}+u_{t}$
where $\theta_{-j}^{*}=(\theta_{k}^{*})_{k\neq j}$ is the vector of coefficients excluding the $j$-th entry, the generic estimator $\widehat{\theta}_{j}^{({\rm lin})}$ can be decomposed into
$\widehat{\theta}_{j}^{({\rm lin})}=\theta_{j}^{*}+\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}\theta_{-j}^{*}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}=:\theta_{j}^{*}+{\rm N}_{j}+{\rm B}_{j},$
@@
For OLS, the score vector $r_{j}=w_{j}^{\perp}$ is orthogonal to the column space of $W_{-j,\cdot}:=(W_{-j,0},\dots,W_{-j,n-1})^{\top}$, under which ${\rm B}_{j}=0$ and no bias is present. The bias term ${\rm B}_{j}$ pops up whenever $r_{j}$ is not orthogonal to $W_{-j,\cdot}$, which happens if we add a penalty to the OLS objective function. With the LASSO penalty at place, we call ${\rm B}_{j}$ shrinkage bias.
Following zhang2014confidence, we replace the unknown parameter $\theta_{-j}^{*}$ by the feasible workhorse estimator $\widehat{\theta}_{-j}^{{\rm S}}=(\widehat{\theta}_{k}^{{\rm S}})_{k\neq j}$ to obtain
@@
where $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ is the Slasso residual. Though LASSO may shrink $\widehat{\theta}_{j}^{{\rm S}}$ all the way to exactly zero, the second term in (2.6) is continuously distributed and therefore $\widehat{\theta}_{j}$ is a desparsified version of $\widehat{\theta}_{j}^{{\rm S}}$. A straightforward calculation yields
$\widehat{\theta}_{j}-\theta_{j}^{*}=\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j}^{{\rm S}})}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}={\rm N}_{j}-(\widehat{{\rm B}}_{j}-{\rm B}_{j}),$ (2.7)
@@
Let $\omega_{j}$ denote the standard deviation of ${\rm N}_{j}$. To secure asymptotic normality for $\widehat{\theta}_{j}$, we need an appropriate score vector $r_{j}$ such that as $n\to\infty$:
@@
The result (R1) is well established by zhang2014confidence under i.i.d. data when the score $r_{j,t}$ is taken as the LASSO residual of regressing $w_{j,t}$ on all other regressors. However, when $w_{j,t}$ is an LUR, this practice leads to a highly persistent $r_{j,t}$ and ruins the asymptotic normality of ${\rm N}_{j}$. This is the Stambaugh bias — a highly persistent regressor produces an asymptotically non-normally distributed OLS estimator skewed away from zero. To safeguard (R1), we must seek a score that is less persistent than an LUR.
@@
${\rm\widehat{B}}_{j}-{\rm B}_{j}$ $=g_{j}^{\top}(\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}),\ \text{where}\ g_{j}=\dfrac{\sum_{t=1}^{n}W_{-j,t-1}r_{j,t-1}}{\sum_{t=1}^{n}w_{j,t-1}r_{j,t-1}}\cdot$ (2.8)
Note that $|{\rm\widehat{B}}_{j}-{\rm B}_{j}|\leq\|g_{j}\|_{\infty}\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}.$ Slasso’s $L_{1}$ estimation error $\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}$ is invariant with the choice of score vector $r_{j,t}$, and diminishes if $\widehat{\theta}^{{\rm S}}$ is consistent. Thus, it suffices to control the order of $\|g_{j}\|_{\infty}$ to achieve (R2). The explicit expression of $g_{j}$ in (2.8) suggests that a weak correlation in sup-norm between $W_{-j,t}$ and the score $r_{j,t}$ (relative to the correlation between $w_{j,t}$ and $r_{j,t}$) helps.
With these routes in mind, we devise XDlasso in the following section. It proceeds with two key steps: (i) conducting an IVX transformation to get a new variable less persistent than the LUR regressors to eliminate the Stambaugh bias; (ii) running an auxiliary LASSO regression to construct an $r_{j,t}$ to remove the shrinkage bias.
@@ 2.3 IVX-Desparsified LASSO
For inference of $\theta_{j}^{*}$ in low dimensions, phillips2009econometric’s IVX method generates an instrument by quasi-differencing $w_{j,t}$. When $w_{j,t}$ is an LUR, this self-generated instrument is mildly integrated. The mitigation of persistence will remove the Stambaugh bias when the sample size passes to infinity, and thus recovers asymptotic normality of the IVX estimator. Due to the coexistence of Stambaugh bias and shrinkage bias, the wisdom of IVX cannot be directly transplanted into the high dimensional case. Instead, we integrate IVX with the idea of desparsified LASSO after comprehending the mechanism of the biases illustrated in Section 2.2.
@@
where $\rho_{\zeta}\in(0,1)$ is a user-determined tuning parameter. Define the s.d. of the instrument as $\widehat{\varsigma}_{j}=\sqrt{n^{-1}\sum_{t=0}^{n-1}\left(\zeta_{j,t}-\bar{\zeta}_{j}\right)^{2}.}$ We unify the scale for LURs and stationary regressors by standardizing the instrument with its s.d.:
$\widetilde{\zeta}_{j,t}=\zeta_{j,t}/\widehat{\varsigma}_{j}.$ (2.10)
@@
For low dimensional predictive regressions, the IVX literature has established asymptotic normality of ${\rm N}_{j}$ in (2.7) taking $r_{j,t}=\widetilde{\zeta}_{j,t}$; see phillips2009econometric. When $w_{j,t}$ is an LUR, the IV $\zeta_{j,t}$ is mildly integrated and less persistent than an LUR. Furthermore, recall that the vector of regressors $W_{-j,t}$ includes either LURs or stationary regressors. Due to different degrees of persistence, the mildly integrated IV $\widetilde{\zeta}_{j,t}$ and the regressors $W_{-j,t}$ are asymptotically uncorrelated. Thus, when $w_{j,t}$ is an LUR, we can choose the score vector as $r_{j,t}=\widetilde{\zeta}_{j,t}$ to deliver a small order of $\|g_{j}\|_{\infty}$. If $w_{j,t}$ is stationary, however, this score vector fails. When $n$ is large, the instrumental variable $\zeta_{j,t}$ behaves similarly as the stationary regressor $w_{j,t}$, and thus its correlation to the high dimensional stationary regressors in $W_{-j,t}$ is not negligible.
@@
The above analysis implies that we must decorrelate the score vector with the other regressors $W_{-j,t}$ to reduce the order of $g_{j}$ to control for the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$. This decorrelation will produce a unified testing approach for both LURs and stationary regressors. To this end, we construct a residual score vector $\widehat{r}_{j}=(\widehat{r}_{j,0},\cdots,\widehat{r}_{j,n-1})^{\top}$ by the following auxiliary LASSO regression
$\widehat{r}_{j,t}$ $=\widetilde{\zeta}_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)},\ \text{where}$ (2.11)
$\widehat{\varphi}^{(j)}$ $=\arg\min_{\varphi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{\zeta}_{j,t-1}-W_{-j,t-1}^{\top}\varphi)^{2}+\mu_{j}\|D_{-j}\varphi\|_{1}$ (2.12)
@@
with the LASSO tuning parameter $\mu_{j}$ and $D_{-j}={\rm diag}(\{\widehat{\sigma}_{k}\}_{k\neq j})$. In low dimensional multivariate predictive regressions, IVX transforms each regressor into a less persistent instrumental variable parallel to (2.9), and constructs a two-stage least squares estimator using all the self-generated instrumental variables. In contrast, we only transform the variable of interest $w_{j,t}$ and estimate one auxiliary regression (2.12).
@@
The score vector $\widehat{r}_{j}$ in (2.11) accommodates stationary and LUR regressors. Recall that when $w_{j,t}$ is stationary, the magnitude of the instrument $\zeta{}_{j,t}$ behaves similarly as $w_{j,t}$. Thus, the score $\widehat{r}_{j}$ is asymptotically equivalent to the standardized residual of the LASSO regression of $w_{j,t}$ on $W_{-j,t}$. The latter is proportional to the score in zhang2014confidence for cross-sectional data, which is also used in adamek2022lasso for stationary time series. When $w_{j,t}$ is an LUR, the instrument $\zeta_{j,t}$ is mildly integrated and has a different degree of persistence from $W_{-j,t}$. Therefore, $\zeta_{j,t}$ is asymptotically uncorrelated to each regressor in $W_{-j,t}$. For notational conciseness, define $\widetilde{W}_{-j,t}=(D_{-j})^{-1}W_{-j,t}$ and $\widetilde{\varphi}^{(j)}=D_{-j}\widehat{\varphi}^{(j)}$. The analysis above suggests $\widetilde{\varphi}^{(j)}\approx 0$ so that $\widehat{r}_{j,t}=\widetilde{\zeta}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\varphi}^{(j)}\approx\widetilde{\zeta}_{j,t}$. Recall from the discussion right after (2.10) that the standardized instrument $\widetilde{\zeta}_{j,t}$ is a valid score process for an LUR regressor, and thus the score $\widehat{r}_{j,t}$ can also remove the biases asymptotically. As a result, the residual of the auxiliary LASSO regression provides a unified construction of the score for either stationary or nonstationary $w_{j,t}$. It allows practitioners to conduct hypothesis testing on the coefficient in high dimensional predictive regression regardless of the order of integration of $w_{j,t}$. We maintain an agnostic attitude about the persistence of the regressors; in practical implementation, we need not distinguish these two types.